# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 21 Apr 2026 v4

@@ 8 Concluding Remarks
remains open; by Proposition 6, the latter would give an independent Fourier-analytic proof of pure singularity. The numerical evidence of Remark 2 and Figure 3 suggests slow decay, but the oscillatory structure makes rigorous conclusions difficult. Determining a global modulus of continuity of $G$ is also open. Table 3 shows $1/c(p)\leq 4$ for all surveyed moduli; Corollary 8 gives $1/c(p)=O(\log\log p)$ pointwise. Closing the gap between the empirical bound $4$ and the asymptotic order $\log\log p$ remains an interesting open problem.
Declaration of competing interests
The author declares that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Research Data Statement
No external datasets were used; all numerical values are reproducible from the explicit formulas and Monte Carlo procedures described in the paper. All empirical values of $c(p)$ in Figures 12 and Table 3 were computed by evaluating the totient ratio $\varphi(p-1)/(p-1)$ from the factorization of p1p-1, followed by the truncated Euler product $P_{J}(p)$ of Proposition 8 with $J=20$; the resulting error is below $10^{-9}$ for every prime $p\geq 3$. The Monte Carlo approximation of $G$ in Figure 1 samples the product distribution of Lemma 10 truncated at the 10,00010{,}000th odd prime $\ell_{N}$. Write $T_{N}:=\sum_{\ell>\ell_{N}}Z_{\ell}$ for the omitted tail and $S_{N}:=\sum_{3\leq\ell\leq\ell_{N}}Z_{\ell}$ for the retained partial sum, so that $X_{N}:=\tfrac{1}{2}e^{-S_{N}}$ approximates $X=\tfrac{1}{2}e^{-S}$ with $S=S_{N}+T_{N}$. The tail mean satisfies $\mathbb{E}[T_{N}]=\sum_{\ell>\ell_{N}}\frac{1}{\ell-1}\log\frac{\ell}{\ell-1}<10^{-5}$. Since $T_{N}\geq 0$ and $X_{N}\leq\tfrac{1}{2}$, the pointwise coupling bound $0\leq X_{N}-X=X_{N}(1-e^{-T_{N}})\leq T_{N}/2$ holds almost surely, giving $W_{1}(G_{N},G)\leq\tfrac{1}{2}\,\mathbb{E}[T_{N}]<5\times 10^{-6}$ for the Wasserstein-$1$ distance between the truncated and full distribution functions. Markov’s inequality further gives $\mathbb{P}(T_{N}>\varepsilon)<10^{-5}/\varepsilon$ for every $\varepsilon>0$. Figure 3 was produced by evaluating the closed-form Euler product of Proposition 5 over all primes $\ell\leq 10^{7}$. Tables 1 and 2 were computed from the explicit formula $B(K)=P_{3}\prod_{i=1}^{K}(1-1/p_{i})$ of Theorem 9 and the exact factorizations of p1p-1 for the listed moduli, respectively, using only rational arithmetic.

### 06 May 2026 v5

@@ 9. Concluding Remarks and Open Questions
remains open. If $\widehat{\mu_{f}}(\tau)\to 0$, the measure $\mu_{f}$ would be a Rajchman measure, placing the shifted-prime distribution of $\log(n/\varphi(n))$ in the same class as many classical singular measures arising from Bernoulli convolutions. A positive $\limsup$ would instead give an independent Fourier-analytic proof of pure singularity via the Riemann–Lebesgue lemma. Numerical evidence (Remark 6.8) suggests slow decay, but the oscillatory structure—driven by the mutual incommensurability of the values $\log(\ell/(\ell-1))$ across primes $\ell$—makes rigorous conclusions difficult. Standard techniques for Bernoulli convolutions (see, e.g., Peres, Schlag, and Solomyak [26]) do not apply directly, since the contraction ratios vary with $\ell$ and do not arise from a single algebraic parameter.
2. Global modulus of continuity. Proposition 6.9 determines the behaviour of $G$ near its right endpoint: $1-G(1/2-\varepsilon)\sim\kappa/\!\log(1/\varepsilon)$, and Corollary 6.11 shows that $G$ is not Hölder of any positive order at $1/2$. Whether a logarithmic modulus of continuity holds globally—that is, whether there exists a constant $C>0$ such that $|G(\alpha)-G(\beta)|\leq C/\!\log(1/|\alpha-\beta|)$ for all $\alpha,\beta\in[0,\tfrac{1}{2}]$ with $|\alpha-\beta|$ small—remains open. For the analogous distribution of $\varphi(n)/n$ over the integers, the modulus of continuity is likewise unknown; see Schoenberg [30], Deshouillers–Hassani [7], and Toulmonde [35] for partial results in the integer case.
Data availability statement
All numerical computations are fully described by the explicit formulae in the paper and are reproducible from these descriptions alone—see Remark 5.15 for truncation and sampling parameters, and Remark 6.8 for the Fourier-decay evaluation. Python scripts reproducing Figures 13 and independently verifying Tables 13 are publicly available at https://github.com/Crypto-Vipin/primitive-root-ratio. No experimental data beyond the computations described in the text were used.